RC meeting and Run Plan, Friday November 25th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information


RC Updates


  • Finished the beam set up, thanks Jiwan and Bill.
  • Running on H2 during Swing and Owl shifts.
  • Started pion run after a run with empty cell.
    • DAC was changed to -120mV
    • We are running on H2, maybe we will run with N2, and we are negotiating to get D2.
  • SBS GEMs are integrated to the DAQ (thank you Xinzhan for the fast delivery)

Short term run plan

  • Sieve slit installation on Saturday and then optics.
  • Getting deuterium to be delivered into the hall tomorrow. (Dave Meekins is the only authorized to deliver Deuterium due to DOE regulations, and he is not available next week).
  • Prepare to install another cell (target group)
  • Todd will be needed to connect the bottle to the reference cell on Monday.


  • Hodoscope TDC was showing left side noisier than right. Contacted Rachel about it and we agreed to investigate first with previous replays and possibly reboot the ROC when possible.
  • Dump right Ion chamber cabling needs to be changed (Jerry Kowal is aware of it)

Hall Access

  • Target group to repair the cell
    • Notice the we will be in Controlled Access until we have full authorization from Radcon and Arun.
  • Ralph to repair or investigate dead hodoscope channels. Ladder to access the BB platform.
  • Todd will connect the deuterium bottle to the reference cell.

Summary of GEn data taking

As soon I get where to get the info, I will update this part.